OpenStudy (anonymous):

Indicate which property is illustrated in Step 3. Step 1 4x + 2 + 8x + 2 = (4x + 2) + (8x + 2) Step 2 = 4x + (2 + 8x) + 2 Step 3 = 4x + (8x + 2) + 2 Step 4 = (4x + 8x) + (2 + 2) Step 5 = (4 + 8)x + (2 + 2) Step 6 = 12x + 4 commutative distributive arithmetic fact associative

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

commutative: a + b + c = a + c +b but also includes: a + (b + c) = a + (c + b)

OpenStudy (anonymous):

would this still be commutative or arithmetic fact @jigglypuff314 becuase im confused about this one abit

OpenStudy (anonymous):

ok i think im understanding it a bit more